Kimberly is twice as old as Ryan. In 4 years, she will be 7 years older than Ryan. How old is Kimberly now?
Answer by ankor@dixie-net.com(22740)

Let k = Kimberly's present age
Let r = Ryan's present age
:
Write an equation for each statement:
:
"Kimberly is twice as old as Ryan."
k = 2r
:
" In 4 years, she will be 7 years older than Ryan."
k + 4 = (r + 4) + 7
k + 4 = r + 11
k = r + 11 - 4
k = r + 7
:
How old is Kimberly now?
:
Replace k with 2r in the above equation
2r = r + 7
2r - r = 7
r = 7 is Ryan's present age
then
2(7) = 14 is K's present age
:
:
Check solution in the statement:
"In 4 years, she will be 7 years older than Ryan."
14 + 4 = (7 + 4) + 7